OpenStudy (anonymous):

Choose one of the factors for 1331x^3-8y^3 a.4 b.11x-2y c.121x^2+44xy+4y^2 d.121x^2-22xy+4y^2

OpenStudy (ivettef365):

1331x^3 = 11 x 11 x 11 x X x X x X and 8y^3 = 2 x 2 x 2 x Y x Y x Y so you have 11x - 2 y

OpenStudy (ivettef365):

Ax^3-Bx^3) = (A-B)(A^2-AB +B^2)